What is the value of the given expression:
log sq.rt.(m^3/n)
a. 1/2 log m - log n
b. 3/2 log m - log n
c. (3 log m - log n)/2
d. (log m - log n)/2

Answer by Jk22(389)

log sq.rt.(m^3/n)=1/2*log(m^3/n)=1/2*(3log(m)-log(n))

->c.